Grade 11General Physics

If the error in measurements of velocity of a particle is 10% and mass is known exactly. The percentage error in K.E. Is?

Ajay Mittur
10 Years agoGrade 11

1 Answer

Aryan Gupta
10 Years ago
If you mean that the mass measured was 100% precise (i.e. no error or 0% error),
the percentage error in the K.E. is 20%.

Solution:
 
error in mass = 0%
error in velocity = 10%
K.E. = m*v*v
therefore % error in K.E. = 0+10+10= 20%.
 
regards
and please approve my answer!